Fermi National Laboratory
 

Operation Summary for February 2, 2011 - February 4, 2011

 

Wednesday Feb. 2

The day shift began with Operations monitoring the Tevatron (TEV) store 8467 (51.47E30), with the Recycler stash lost due to a breaker trip, with Operations monitoring the Antiproton Source (Pbar) stack (11.9E10), with the Main Injector (MI) off due to the breaker trip, with Linac and Booster operating normally, with the NuMI, MiniBooNE and Meson FTBF T-978 (CALICE) experiments off, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.

During the day shift, Controls personnel traced a problem with a MiniBooNE beam detector to bad AART cards and motherboards; they had them fixed by 11:45 AM. Operations reestablished beam to MiniBooNE at 12:07 PM. High Voltage personnel investigated high voltage system problems all day: they reported two power lines were down in the Village as well as the breaker trip at the Kautz Rd. substation. At the end of shift, FESS personnel were working on the MI-10 transformer and retrieving test equipment from the Master Substation. Operations reported a short downtime due to a quadrupole power supply (E:Q847).

During the evening shift, Operations reported that MI was still off, but that operators had established beam to MiniBooNE. Operations reported that the MI-10 fault cleared after personnel cycled the L10 switch-gear. Operators turned on MI. Operations reported that a Linac RF station (LRF3) tripped off; Linac personnel came in and tuned the station. A Pbar critical device failed to open; Interlocks personnel came in and determined that a controller card had failed. Operators reported that VDPA1 and 2 power supplies had LCW leaks; EE Support personnel came in and tightened hoses. MI-10 injection kickers KPS1A, B, and C would not reset; Kicker personnel came in and found a problem with the RF signal going to MI-10 from the Booster LLRF room.

Thursday Feb. 3

The midnight shift began with Operations monitoring store 8467 (25.78E30), with no stash and with operators preparing the Recycler for an access, with Operations monitoring the Pbar stack (11.8E10), and with NuMI and MiniBooNE off.

During the midnight shift, Kicker personnel accessed MI to fill the Recycler kicker magnet with Fluorinert; they checked and filled the other kicker magnets at the same time. Kicker personnel reported that a RF splitter had caused the MI-10 Kicker RF signal issue; they replaced the splitter. Operations had beam to MiniBooNE by 3:32 AM. A NuMI power supply (V108) would not come on; this held off beam to NuMI. EE Support technicians came in to make repairs. Operations established beam to Pbar at 4:28 AM, followed shortly by stashing to the Recycler. Operations established beam to T-978 at 5:13 AM. Operations reported that EE Support techs had completed repairs on V108 at 6:13 AM and NUMI resumed taking beam.

During the day shift, at 10:07 AM, Operations reported that the stash was lost when breaker 87-1 tripped at the Kautz Rd. substation and MI went unsecure from MI-40 to MI-10. This trip halted beam to Pbar, NuMI, MiniBooNE, and T-978. FESS personnel traced the breaker trip to the MI-10 quadrupole power supply transformer on the primary side, which had arced over due to a buildup of ice in the breaker cabinet. TEV personnel conducted Tevatron Electron Lens studies from 9:33 AM until the store aborted at 12:51 PM due to bad beam positions at TEV sector E1. Operators prepared CDF, D0, MI and Booster for accesses. By the end of shift, both Booster and the TEV turn-on were in progress and the MI access was finishing up.

During the evening shift, Booster turn-on continued. Operators rebooted MECAR and MIBEAM. Booster and MI resumed sending beam at 5:41 PM. Operations reported that a quadrupole power supply (I:Q713) held off stacking and it needed a local reset. Pbar resumed stacking at 6:17 PM. CDF made a 30-minute access. Afterwards, the TEV turn-on ran until the end of shift. A Linac Klystron (KRF6) tripped on a transformer LCW flow fault; Water techs came in and replaced the flow meter. Operations reported that NuMI had been off all shift due to a VCB turn on fault; power supply techs worked on the problem for several hours; however, they were unable to repair the supply; they decided to suspend repairs until Friday morning due to the frigid weather conditions. Operations delivered beam to MiniBooNE when available.

Friday Feb. 4

The midnight shift began with Operations monitoring the stash (103E10), with Pbar stacking (25.5E10), with NuMI off, and with MiniBooNE taking beam.

During the midnight shift, operators completed a series of TEV wet squeezes to check and adjust orbits. Operations had the TEV in shot setup by 2:01 PM and established store 8473 at 3:25 AM with an average initial luminosity of 93.14E30. Operations reported that problems with the Recycler dampers during shot setup caused the loss of ~50E10 antiprotons, which caused the low luminosity. Operators called in Damper personnel to troubleshoot the system. Antiprotons were stacked and transferred to the Recycler as appropriate. Operations established beam to T-978 at 4:42 AM.

The Plans for Friday and the Weekend

The plans for today are to get NuMI back online and to stack, stash, and store for the rest of the weekend.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 02/04/2011    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory